# SpeechAct: Towards Generating Whole-body Motion from Speech

### v3

@@ Abstract
Whole-body motion generation from speech is crucial for computer graphics and immersive VR/AR. Prior methods struggle to produce reasonable and diverse whole-body motions from speech. In this paper, we introduce a novel method, named SpeechAct, based on a hybrid point representation and contrastive motion learning to boost realism and diversity in motion generation. Our hybrid point representation leverages the advantages of keypoint representation and surface points of 3D body model, which is easy to learn and helps to achieve smooth and reasonable motion generation from speech. We design a VQ-VAE to learn a motion codebook using our hybrid presentation, and then regress the motion from the input audio using a translation model. To boost diversity in motion generation, we propose a contrastive motion learning method according to the intuitive idea that the generated motion should be different from the motions of other audios and other speakers. We collect negative samples from other audio inputs and other speakers using our translation model. With these negative samples, we pull the current motion away from them using a contrastive loss to produce more distinctive representations. In addition, we compose a face generator to generate deterministic face motion due to the strong connection between the face movements and the speech. Experimental results validate the superior performance of our model. The code will be available for research purposes.
@@ 1 Introduction
This file is intended to serve Human motion generation from speech is a critical area in computer graphics and immersive VR/AR [1, 2, 3, 4], which has been extensively studied as a way of human behavior understanding [5]. Given a speech recording, the goal is to generate a spectrum of diverse, yet reasonable motion sequences, which is in line with real-life scenarios and can meet the varying needs of different individuals.
@@
Existing approaches concentrate on translating speech to the motion of the partial body [6, 7, 8, 9]. They use keypoints of the body as their motion representation, which is easy to learn and contains local details for hands. However, the keypoints leave degrees of freedom undefined, e.g., the rotation of a limb, leading to inaccurate and unrealistic results when fitting or animating a full 3D body. For the whole-body motion generation, TalkShow [5] adopts the parametric representation of SMPL-X body mesh as the motion representation, and proposes a two-stage model to recover the full 3D body. However, we observe that the motion generated by TalkShow can generate discontinuous results, for example, jitters and disharmony contact between the feet and the ground. This is because they use parametric representation defined in the axis-angle space using the kinematic tree, which may introduce the discontinuity caused by the axis-angle representation, leading to the difficulty of modeling continuous latent space [10, 11, 12]. Moreover, due to the inherent one-to-one mapping between input and output during the training process, all the above approaches can tend to learn an average motion lacking diversity. However, in fact, given a speech recording, different speakers tend to exhibit varying motions in different situations. Therefore, how to generate diverse results is also a challenging problem.
@@
In this work, to generate smooth and reasonable motion, we propose a hybrid point representation for whole-body motion generation from speech. Specifically, the keypoint representation is easy to learn and contains local details for hands, but lacks surface constraints to obtain the 3D body model. To combat this problem, we adopt the surface points of SMPL-X body mesh as a part of our hybrid point representation to eliminate the ambiguity due to the keypoint representation. Therefore, defined in Euclidean space, our hybrid point representation is easy to learn and encompasses global constraints and local details for whole-body motion generation. This helps generate smooth and reasonable results, e.g., avoiding foot skating. Besides, to provide an easy-to-use output representation that can be used in many applications, i.e., SMPL-X body mesh, we design a generator to recover the parameters of SMPL-X from our hybrid representation.
@@
With this hybrid point representation, we can generate smooth and reasonable results, however, similar to previous approaches, the diversity of generated results is still limited due to the one-to-one mapping of the training data. This one-to-one mapping makes the model capture similar motions, e.g., certain habitual body motions of a speaker and similar body motions of different speakers due to specific meanings, but this limits the model to generate diverse motions. To boost the diversity of generated results, an intuitive idea is that, the generated motion of a specific audio should be different from the generated motions of other audios and other speakers. To achieve this, we introduce a novel contrastive motion learning method to obtain distinctive motion representations. Specifically, we first collect negative samples that are discrete motion representations from other audio inputs and other speakers using the translation model. Then, we pull the current generated motion away from the negative samples by a contrastive loss, and thus the generated motion can be more distinctive. By using the target motion as a positive sample, our approach encourages diverse motion learning while capturing similar motions.
@@
Built on our hybrid point representation and motion contrastive learning method, we present a novel model, named SpeechAct, to generate reasonable and diverse body motion, which first constructs a quantized motion codebook based on our presentation, and then regresses the distinct motion from speech using our contrastive motion learning. In addition, we compose a face generator to generate deterministic face motion due to the strong connection between the face movements and the speech. Experimental results compared with several state-of-the-art methods validate the effectiveness of our model. We give an application that animates avatars using speech inputs, which can be further used in immersive AR/VR. Moreover, our model can generate promising results conditioned on different languages and music. Fig. 1 presents two samples generated by our model. The human body meshes corresponding to the same text color indicate the motion generated by the driven speech content. It can be seen that our approach not only generates smooth sequences of movements containing diverse poses, but also ensures that these sequences correspond to the audio content.
@@
  • We design SpeechAct, a novel framework to generate whole-body motion from speech, which can produce more reasonable and diverse results.
@@
  • Experimental results demonstrate our model can generate natural and diverse results. Our method can also be generalized to other languages and music inputs.
@@ 2.1 Human Motion Generation
Human motion generation is the task of synthesizing realistic and natural human movements for various applications, such as animation, gaming, robotics, and VR/AR. The data representations can be divided into two kinds, i.e., keypoint-based and rotation-based. The keypoint-based representation [7, 13, 14] adopts a set of keypoints that are specific points on the body, e.g., joints. The keypoints are defined using 2D/3D coordinates in the pixel or world coordinate system. Some works [12, 15] use sparse markers to represent human pose, which can provide more comprehensive information for rotation. However, the keypoint-based methods can provide local details for hands, but can not constrain the global shape, such as the rotation of limbs, which can lead to unrealistic results when fitting or animating a full 3D body. The rotation-based methods [16, 17] adopt the rotation of the body part relative to their parent, which helps to constrain the body pose by modeling the body, face and hands jointly. However, this representation can generate discontinuous results due to its specific space [10, 11, 12], e.g., axis-angle space for SMPL-X body model.
@@ 2.2 Whole-body Motion Generation from Speech
Some approaches cooperate GANs [29, 30, 31, 4], VAEs [32, 3, 33], VQ-VAEs [8, 34], and normalizing flows [35, 36] to increase the diversity of the generated results, but the results are inadequate [5]. Recently, due to the great success of diffusion models [37], some methods [38, 39, 40] introduce diffusion models to generate human motion from speech, which can generate promising results. However, this kind of method requires a large number of diffusion steps, and thus has low generation efficiency. Besides, modeling part of the whole-body is insufficient for a comprehensive understanding of human behavior. To generate whole-body motion, most similar to our work, Yi et al. [5] fit the SMPL-X body mesh based on the speaker-specific gesture dataset [41], but the motions in their dataset focus on face and upper-body gestures. Based on their dataset, they propose to generate diverse motions for the body and hands using two independent codebooks. Liu et al. [42] propose a new whole-body co-speech dataset, i.e., BETAX dataset, and design a masked audio gesture transformer to regress the gesture from the audio input. However, because of the inherent one-to-one mapping between input and output, the model struggles to generate reasonable and diverse results. Besides, they adopt parameters of SMPL-X body mesh as the motion representation, which introduces complexity for the prediction task [43] and can lead to discontinuous problems.
@@ 3 Method
Given a speech recording, our work aims to generate diverse whole-body motion sequences that are in harmony with the provided speech. We design a novel framework, named SpeechAct, to achieve this. The most significant differences with existing approaches are that, to generate reasonable and diverse results for the body and hands, we propose a hybrid point representation to form our motion efficiently, and a contrastive motion learning method to distinguish the generated motion from other motions to boost the diversity of the generated results. Fig. 2 shows the overall framework of our SpeechAct. Our model consists of two generators: a body generator (Sec. 3.1) and a face generator (Sec. 3.2). Because the audio signal is not very closely related to hand and body gestures [5], the body generator aims to generate reasonable and diverse results. For an effective and versatile motion representation, we introduce a hybrid point representation (Sec. 3.1.1) and develop a generator to obtain SMPL-X model parameters. To ensure the generation of coherent and diverse results, we leverage advancements in VQ-VAE to learn a quantized motion space (Sec. 3.1.2). Building upon this quantized motion space, we propose a translation model (Sec. 3.1.3) and introduce a novel contrastive motion learning method, allowing us to transform audio signals into motion representations while promoting diversity in body and hand motions. In contrast, speech content, especially phonetic information, is closely related to face motion, notably lip motion, thus the face generator is designed using an encoder-decoder architecture and is responsible for generating deterministic results.
@@ 3.1 Body Generator
For the body generator, we aim to generate reasonable and diverse motion sequences for the body and hands given an audio input. Previous approaches [6, 7] adopt keypoints for motion representation, limiting their ability to fully reconstruct or animate a 3D body. TalkShow [5] utilizes SMPL-X model parameters, resulting in discontinuous motion sequences. To generate more realistic and continuous results, we propose a hybrid point representation with a generator, which can generate continuous results efficiently and deliver the SMPL-X model accurately. Besides, because the input and the output are one-to-one mappings, previous methods struggle to generate reasonable yet diverse results due to the challenges of balancing constraints between precision and diversity. To alleviate this problem, we design a two-stage model and introduce a novel contrastive motion learning method to boost the diversity of generated results. Fig. 2 shows the framework of our body generator. Specifically, we first learn a motion codebook by encoding motion sequences into quantized codes. To synthesize more natural hand motion, we model the body and hands separately using three different codebooks. Subsequently, our translation model autoregressively transforms audio signals into motion sequences. With the contrastive motion learning method, the translation model can generate more distinctive motion representations, which helps to generate diverse results.
@@ 3.1.1 Hybrid Point Representation
As shown in the top-left corner of Fig. 2, our hybrid points representation consists of two vital components: surface points $v_{s}$ (the points in light blue) and keypoints $v_{k}$ (the points in dark blue). Surface points are extracted from the SMPL-X body mesh following [49], serving as a global constraint to the freedom of the body. Keypoints, on the other hand, capture fine-grained details, particularly in areas like the hands, enriching the representation. The final representation $v$, results from the concatenation of $v_{s}$ and $v_{k}$, with each element representing a 3D location. This combined representation facilitates the capture of both global constraints and local details. This representation is employed to represent a motion sequence, denoted as $V=\{v_{t}\}_{t=0}^{T}$. It encapsulates the dynamic evolution of the hybrid point-based representation over time, enabling comprehensive modeling and analysis of human body motion.
Inspired by [10], to eliminate the gap between the SMPL-X body mesh and our representation, we design a simple generator $\mathcal{G}_{P}$, where the input is $V$ and the output is the parameters of the SMPL-X body mesh. This generator $\mathcal{G}_{P}$ consists of a series of neural network layers, specifically a stack of three residual blocks, each with varying input and output channels. These layers are responsible for transforming the input point representation into a more meaningful intermediate feature representation. Subsequently, a final convolutional layer is employed to produce the desired output, which represents the parameters of the SMPL-X body mesh. By doing so, even training the body and face separately, our model can ensure consistency between body and face to generate seamless and continuous results. This generator is trained with our body generator, which is represented in Sec. 3.1.2.
@@ 3.1.2 Quantized Motion Modeling
As shown in Fig. 2 (a), the first stage of our body generator is quantized motion modeling. Given a set of motion sequences $M=\{m_{t}\}_{t=0}^{T}$, we first extract our hybrid point representation $V=\{v_{t}\}_{t=0}^{T}$. With our hybrid point representation, we adopt a Vector Quantized-VAE (VQ-VAE) [50] to learn a meaningful and compact motion space. We first obtain motion features $Z\in\mathcal{R}^{\frac{T}{w_{m}}\times n_{z}}$ from the input motion sequence, where $n_{z}$ is the channel dimension of quantized features and $w_{m}$ is the temporal window size.
Denote the learneable motion codebook as $C=\{c_{i}\}_{i=1}^{N}$, where $N$ is the length of the codebook. Then, we quantize the motion features by replacing them with the nearest codes in the codebook. Specifically, for each feature $z_{t}\in Z$, the quantized feature can be obtained by:

### v5

@@ Abstract
Whole-body motion generation from speech audio is crucial for computer graphics and immersive VR/AR. Prior methods struggle to produce natural and diverse whole-body motions from speech. In this paper, we introduce a novel method, named SpeechAct, based on a hybrid point representation and contrastive motion learning to boost realism and diversity in motion generation. Our hybrid point representation leverages the advantages of keypoint representation and surface points of 3D body model, which is easy to learn and helps to achieve smooth and natural motion generation from speech audio. We design a VQ-VAE to learn a motion codebook using our hybrid presentation, and then regress the motion from the input audio using a translation model. To boost diversity in motion generation, we propose a contrastive motion learning method according to the intuitive idea that the generated motion should be different from the motions of other audios and other speakers. We collect negative samples from other audio inputs and other speakers using our translation model. With these negative samples, we pull the current motion away from them using a contrastive loss to produce more distinctive representations. In addition, we compose a face generator to generate deterministic face motion due to the strong connection between the face movements and the speech audio. Experimental results validate the superior performance of our model. The code is available at http://cic.tju.edu.cn/faculty/likun/projects/SpeechAct/index.html
@@ 1 Introduction
Human motion generation from speech audio [1, 2, 3] is a critical area in computer graphics and immersive VR/AR [4, 5, 6, 7], which has been extensively studied as a way of human behavior understanding [1]. Given a speech recording, the goal is to generate a spectrum of diverse yet natural motion sequences, which is in line with real-life scenarios and can meet the varying needs of different individuals.
@@
Existing approaches concentrate on translating speech to the motion of the partial body [8, 3, 9, 10]. They use keypoints of the body as their motion representation, which is easy to learn and contains local details for hands. However, the keypoints leave degrees of freedom undefined, e.g., the rotation of a limb, leading to inaccurate and unrealistic results when fitting or animating a full 3D body. For the whole-body motion generation, TalkShow [1] adopts the parametric representation of SMPL-X body mesh as the motion representation, and proposes a two-stage model to recover the full 3D body. However, we observe that the motion generated by TalkShow often contains discontinuous results, for example, jitters and disharmony contact between the feet and the ground. This is because they use parametric representation defined in the axis-angle space using the kinematic tree, which may introduce the discontinuity caused by the axis-angle representation, leading to the difficulty of modeling continuous latent space [11, 12, 13]. Moreover, due to the inherent one-to-one mapping in our real life and training pairs, all the above approaches tend to learn an average motion lacking diversity. However, in fact, given a speech recording, different speakers tend to exhibit varying motions in different situations. Therefore, how to generate diverse results is also a challenging problem.
@@
In this work, to generate smooth and natural motion, we propose a hybrid point representation for whole-body motion generation from speech audio. Specifically, the keypoint representation is easy to learn and contains local details for hands, but lacks surface constraints to obtain the 3D body model. To combat this problem, we adopt the surface points of SMPL-X body mesh as a part of our hybrid point representation to eliminate the ambiguity due to the keypoint representation. Therefore, defined in Euclidean space, our hybrid point representation is easy to learn and encompasses global constraints and local details for whole-body motion generation. This helps generate smooth and natural results, e.g., avoiding foot skating. Besides, to provide an easy-to-use output representation that can be used in many applications, i.e., SMPL-X body mesh, we design a generator to recover the parameters of SMPL-X from our hybrid representation.
@@
With this hybrid point representation, we can generate smooth and natural results, however, similar to previous approaches, the diversity of generated results is still limited due to the one-to-one mapping in the training data. This one-to-one mapping in the training data makes the model capture similar motions, e.g., certain habitual body motions of a speaker and similar body motions of different speakers due to specific meanings, but this limits the model to generate diverse motions. To boost the diversity of generated results, an intuitive idea is that, the generated motion of a specific audio should be different from the generated motions of other audios and other speakers. To achieve this, we introduce a novel contrastive motion learning method to obtain distinctive motion representations. Specifically, we first collect negative samples that are discrete motion representations from other audio inputs and other speakers using the translation model. Then, we pull the current generated motion away from the negative samples by a contrastive loss, and thus the generated motion can be more distinctive. By using the target motion as a positive sample, our approach encourages diverse motion learning while capturing similar motions.
@@
Built on our hybrid point representation and motion contrastive learning method, we present a novel model, named SpeechAct, to generate natural and diverse body motion. This model first constructs a quantized motion codebook based on our presentation, and then regresses the distinct motion from speech audio using our contrastive motion learning. In addition, to support whole-body motion synthesis, we compose a face generator to generate deterministic face motion due to the strong connection between the face movements and the speech audio. Experimental results compared with several state-of-the-art methods validate the effectiveness of our model. We give an application that animates avatars using speech inputs, which can be applied in immersive AR/VR [14, 15]. Moreover, our model can generate promising results conditioned on different languages. Fig. 1 presents two samples generated by our model. The human body meshes corresponding to the same text color indicate the motion generated by the driven speech content. It can be seen that our approach not only generates smooth sequences of movements containing diverse poses, but also ensures that these sequences correspond to the audio content.
@@
  • We design SpeechAct, a novel framework based on a new representation to generate whole-body motion from speech audio, which can produce more natural and diverse results.
@@
  • Experimental results demonstrate our model can generate natural and diverse results. Our method can also be generalized to other languages.
@@ 2.1 Human Motion Generation
Human motion generation is the task of synthesizing realistic and natural human movements for various applications, such as animation, gaming, robotics, and VR/AR. The data representations can be divided into two kinds, i.e., keypoint-based and rotation-based. The keypoint-based representation [3, 16, 17] adopts a set of keypoints that are specific points on the body, e.g., joints. The keypoints are defined using 2D/3D coordinates in the pixel or world coordinate system. Some works [13, 18] use sparse markers to represent human pose, which can provide more comprehensive information for rotation. However, the keypoint-based methods can provide local details for hands, but can not constrain the global shape, such as the rotation of limbs, which can lead to unrealistic results when fitting or animating a full 3D body. The rotation-based methods [19, 20] adopt the rotation of the body part relative to their parent, which helps to constrain the body pose by modeling the body, face and hands jointly. However, this representation can generate discontinuous results due to its specific space [11, 12, 13, 21], e.g., axis-angle space for SMPL-X body model.
@@ 2.2 Whole-body Motion Generation from Speech
Some approaches cooperate GANs [33, 34, 35, 7], VAEs [36, 6, 37], VQ-VAEs [9, 38], and normalizing flows [39, 40] to increase the diversity of the generated results, but the results are inadequate [1]. Recently, due to the great success of diffusion models [41], some methods [42, 43, 2, 44] introduce diffusion models to generate human motion from speech, which can generate promising results. However, this kind of method requires a large number of diffusion steps, and thus has low generation efficiency. Besides, modeling part of the whole-body is insufficient for a comprehensive understanding of human behavior. To generate whole-body motion, most similar to our work, Yi et al. [1] fit the SMPL-X body mesh based on the speaker-specific gesture dataset [8], but the motions in their dataset focus on face and upper-body gestures. Based on their dataset, they propose to generate diverse motions for the body and hands using two independent codebooks. Liu et al. [45] propose a new whole-body co-speech dataset, i.e., BETAX dataset, and design a masked audio gesture transformer to regress the gesture from the audio input. However, because of the inherent one-to-one mapping between input and output, the model struggles to generate natural and diverse results. Besides, they adopt parameters of SMPL-X body mesh as the motion representation, which introduces complexity for the prediction task [46] and can lead to discontinuous problems.
@@ 3 Method
Given a speech recording, our work aims to generate diverse whole-body motion sequences that are in harmony with the provided speech. We design a novel framework, named SpeechAct, to achieve this. The most significant differences with existing approaches are that, to generate natural and diverse results for the body and hands, we propose a hybrid point representation to form our motion efficiently, and a contrastive motion learning method to distinguish the generated motion from other motions to boost the diversity of the generated results. Fig. 2 shows the overall framework of our SpeechAct. Our model consists of two generators: a body generator (Sec. 3.1) and a face generator (Sec. 3.2). Because the audio signal is not very closely related to hand and body gestures [1], the body generator aims to generate natural and diverse results. For an effective and versatile motion representation, we introduce a hybrid point representation (Sec. 3.1.1) and develop a generator to obtain SMPL-X model parameters. To ensure the generation of coherent and diverse results, we leverage advancements in VQ-VAE to learn a quantized motion space (Sec. 3.1.2). Building upon this quantized motion space, we propose a translation model (Sec. 3.1.3) and introduce a novel contrastive motion learning method, allowing us to transform audio signals into motion representations while promoting diversity in body and hand motions. In contrast, speech content, especially phonetic information, is closely related to face motion, notably lip motion, thus the face generator is designed using an encoder-decoder architecture and is responsible for generating deterministic results.
@@ 3.1 Body Generator
For the body generator, we aim to generate natural and diverse motion sequences for the body and hands given an audio input. Previous approaches [8, 3] adopt keypoints for motion representation, limiting their ability to fully reconstruct or animate a 3D body. TalkShow [1] utilizes SMPL-X model parameters, resulting in discontinuous motion sequences. To generate more realistic and continuous results, we propose a hybrid point representation with a generator, which can generate continuous results efficiently and deliver the SMPL-X model accurately. Besides, because the input and the output are one-to-one mappings, previous methods struggle to generate natural yet diverse results due to the challenges of balancing constraints between precision and diversity. To alleviate this problem, we design a two-stage model and introduce a novel contrastive motion learning method to boost the diversity of generated results. Fig. 2 shows the framework of our body generator. Specifically, we first learn a motion codebook by encoding motion sequences into quantized codes. To synthesize more natural hand motion, we model the body and hands separately using three different codebooks. Subsequently, our translation model autoregressively transforms audio signals into motion sequences. With the contrastive motion learning method, the translation model can generate more distinctive motion representations, which helps to generate diverse results.
@@ 3.1.1 Hybrid Point Representation
Figure 3: The overview of our hybrid representation
As shown in Fig. 3, our hybrid points representation consists of two vital components: surface points $v_{s}$ (the points in light blue) and keypoints $v_{k}$ (the points in dark blue). Surface points are extracted from the SMPL-X body mesh using mesh sampling operation [52], serving as a global constraint to the freedom of the body. On the other hand, keypoints, which are defined by SMPL-X body mesh and can be obtained using the joint regressor from SMPL-X [53], capture fine-grained details, particularly in areas like the hands, enriching the representation. In practice, we adopt all keypoints, i.e., 55 keypoints, of SMPL-X body mesh, and use 431 points sampled from the surface. As both keypoints and surface points are defined in Euclidean space, the final representation $v\in\mathcal{R}^{486\times 3}$, results from the concatenation of $v_{s}$ and $v_{k}$, with each element representing a 3D location. This combined representation facilitates the capture of both global constraints and local details. This representation is employed to represent a motion sequence, denoted as $V=\{v_{t}\}_{t=0}^{T}$, where $T$ represents the number of frames in a motion sequence. It encapsulates the dynamic evolution of the hybrid point-based representation over time, enabling comprehensive modeling and analysis of human body motion.
Inspired by [11], to eliminate the gap between the SMPL-X body mesh and our representation, we design a simple generator $\mathcal{G}_{P}$, where the input is $v_{t}$ and the output is the parameters of the SMPL-X body mesh. This generator $\mathcal{G}_{P}$ consists of a series of neural network layers, specifically a stack of three residual blocks, each with varying input and output channels. These layers are responsible for transforming the input point representation into a more meaningful intermediate feature representation. Subsequently, a final convolutional layer is employed to produce the desired output, which represents the parameters of the SMPL-X body mesh. By doing so, even training the body and face separately, our model can ensure consistency between body and face to generate seamless and continuous results. This generator is trained with our body generator, which is represented in Sec. 3.1.2.
@@ 3.1.2 Quantized Motion Modeling
As shown in Fig. 2 (a), the first stage of our body generator is quantized motion modeling. Given a set of motion sequences =$\mathcal{M}=\{M_{k}\}_{k=0}^{K}$, where $K$ is the number of motion sequences, we first extract our hybrid point representation $\mathcal{V}=\{V_{k}\}_{k=0}^{K}$. With a sample $V$ from $\mathcal{V}$, we adopt a Vector Quantized-VAE (VQ-VAE) [54] to learn a meaningful and compact motion space. We first obtain motion features $Z\in\mathcal{R}^{\frac{T}{w_{m}}\times n_{z}}$ from the input motion sequence, where $T$ is the number of frames in $V$, $n_{z}$ is the channel dimension of quantized features and $w_{m}$ is the temporal window size.
Denote the learneable motion codebook as $C=\{c_{i}\}_{i=1}^{N}$, where $N$ is the length of the codebook. Then, we quantize the motion features by replacing them with the nearest codes in the codebook. Specifically, for each row $z_{t}$ in $Z$, the quantized feature can be obtained by:
@@
In practice, we adopt three codebooks, i.e., left-hand codebook, right-hand codebook, and body codebook, to model more detailed motions for the body and hands, respectively. Therefore, the encoded feature $Z$ can be split into three parts $Z^{lh},Z^{rh},Z^{b}$ according to different body parts of our hybrid representation before quantization, and these three features can be quantized using Eq. 1. The quantized feature $Z_{q}$ is the concatenation of left-hand quantized feature $Z_{q}^{lh}$, right-hand quantized feature $Z_{q}^{rh}$, and body quantized feature $Z_{q}^{b}$. Therefore, though we adopt three codebooks, they can be optimized using Eq. 2 and Eq. 3 simultaneously.